OpenStudy (alayrhi):

What is −1 2/3⋅(−5 1/4)?

OpenStudy (wholock):

Okay, so is that -12, or -1 * (2/3)?

OpenStudy (alayrhi):

It's -1 and 2/3 times -5 and 1/4

OpenStudy (mathmale):

Convert both mixed numbers into improper fractions. Please show your work. Example: -2 3/7 = -17/7

OpenStudy (wholock):

Okay, well using PEMDAS, you know to solve the equation in the parenthesis first, and so -5 * (1/4) = -5/4, (-1.25 in decimal form) Now, you know that anything times 1 = itself, but in this case you have a negative infront of the one, so you keep the negative giving you: (-1/2) * (-5/4) Plug this into a calculator and you get: 5/8

OpenStudy (alayrhi):

thanks